# See Less, See Right: Bi-directional Perceptual Shaping For Multimodal Reasoning

### v1

@@ 6.1 Data Statistics
  • Stage 1: Sampling and Reformulation. We initially randomly sampled 50,000 raw chart-code pairs from the ECD [yang2025ecd10k] dataset. After the Question Reformulation and Validation phase, where the LLM arbitrator (GPT-5-mini) converted open-ended questions into verified multiple-choice formats, approximately 30K valid samples were retained.
@@ 7.1 Case Study
7.2 Comparison with Standard GRPO
To verify that the improvements stem from our specific perceptual shaping pipeline rather than merely applying RL to the data, we compare BiPS-General against a baseline trained with standard GRPO. This baseline is fine-tuned on the exact same combined dataset (programmatic samples + ViRL39k) but treats all data uniformly without the proposed two-stage perceptual shaping curriculum.
Table 7: Comparison between standard GRPO and our BiPS-General on the same data mixture. Our method consistently outperforms the standard GRPO baseline across all benchmarks.
@@ 7.2 Comparison with Standard GRPO
Qwen2.5-VL-7B 42.5 36.6 26.8 52.0 68.2 25.2 41.1 62.1
GRPO 45.4 50.2 32.9 68.0 74.3 27.3 42.6 64.6
BiPS-General 50.6 51.8 34.0 68.7 75.0 28.6 45.3 65.7
Analysis. As shown in Table 7, while standard GRPO yields substantial gains over the base model, BiPS-General consistently outperforms it across all benchmarks. Notably, on complex chart reasoning tasks like CharXiv, our method surpasses the standard GRPO baseline by a significant margin (+5.2). This performance gap highlights a critical insight: simply optimizing reasoning via RL is insufficient if the underlying visual grounding is flawed. By explicitly shaping the model’s perception through our programmatic curriculum before the general reasoning stage, BiPS ensures that the RL process operates on high-fidelity visual signals, thereby amplifying the effectiveness of the optimization.

### v2

@@ 6.1 Data Statistics
  • Stage 1: Sampling and Reformulation. We initially randomly sampled 50,000 raw chart-code pairs from the ECD [49] dataset. After the Question Reformulation and Validation phase, where the LLM arbitrator (GPT-5-mini) converted open-ended questions into verified multiple-choice formats, approximately 30K valid samples were retained.
@@ 7.1 Case Study
Figure 7: Evidence-Preserving and Evidence-Ablated views.
@@
Figure 7 illustrates an example of an evidence-preserving view and an evidence-ablated view. This case clearly shows that simple operations such as cropping are insufficient for generating the two views, since the relevant visual evidence can be fine-grained and the associated meta-information may be sparse in the image. Moreover, the evidence-preserving view must retain all visual elements related to the question, including the necessary meta-information, whereas the evidence-ablated view provides a more precise, fine-grained modification that removes only the key visual element (e.g., the target line).
7.2 Qwen3-VL-Thinking Results and Discussion
We further evaluate BiPS in the thinking-mode setting by fine-tuning Qwen3-VL-8B-Thinking on 13K chart samples, with hyperparameters following the same configuration as described in Section 4.1 and Section 6. We still obtain significant improvements on charts and generalization to non-chart domains, which shows that BiPS provides perceptual improvements that complement the model’s strong reasoning capabilities.
Setting.
We use temperature = 1.0, top-p = 0.95, top-k = 20 and presence penalty = 1.2 for the fine-tuned model by default. For the base Qwen3-VL-8B-Thinking model, we keep temperature, top-p, and top-k unchanged, and sweep a small, pre-specified grid of presence penalties: 1.2 (matching the fine-tuned setting) and 1.5 (following the technical report [31]). This controlled sweep accounts for decoding sensitivity and maintains consistency with prior evaluations, and we report the best score within this fixed grid for the base model. The prompts follow those used in Qwen3-VL [31]. For MMStar, we remove the duplicate instruction “Please select the correct answer from the options above.” when the question already includes the hint: “Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.” We find that RL fine-tuning on chart samples negatively impacts optical character recognition (OCR) performance on text. This effect can be partially alleviated by adding a system prompt that biases the model toward OCR: “You are a vision-language model. For image-based problems, always prioritize accurate reading of all visible text, symbols, and numbers.” We adopt this prompt for MathVerse, where a subset of examples present the question directly within the image. We use the same prompt across all models for a fair comparison.
Table 7: Effect of BiPS on Qwen3-VL-8B-Thinking
@@ 7.2 Qwen3-VL-Thinking Results and Discussion
Qwen3-VL-8B-Thinking 53.0 54.1 40.4 72.2 81.0 62.2 77.2 75.3
GRPO 54.3 54.6 43.0 74.0 80.4 60.8 76.2 75.3
Ours 58.1 56.8 44.1 75.5 80.4 63.9 77.4 76.3
Discussion.
While BiPS delivers consistently strong in-domain gains and robust OOD generalization on Qwen2.5-VL-7B, its behavior on the more advanced reasoning model Qwen3-VL-8B-Thinking reflects a more challenging generalization setting. BiPS continues to yield substantial in-domain improvements, while largely maintaining cross-domain performance with moderate OOD gains on most benchmarks. This trend is expected: BiPS is trained exclusively on chart data, and transferring fine-grained perceptual supervision to diverse visual domains becomes inherently harder under strong reasoning priors. Nevertheless, BiPS preserves cross-domain performance, indicating that the learned visual grounding signal remains transferable, although its magnitude is naturally bounded when the target domains differ substantially from the training distribution.
Extensibility.
Beyond chart-centric training, BiPS naturally generalizes as a transferable perceptual supervision mechanism. While chart data alone already yields consistent cross-domain improvements, stronger and more uniform gains can be expected by extending BiPS to multi-domain training, for example by 1) mixing heterogeneous data, where non-chart domains are optimized with standard GRPO objectives, or 2) constructing bidirectional views across multiple domains and jointly optimizing them under the BiPS framework. BiPS can be extended to non-chart domains through construction of training views. For example, for natural images, recent visual chain-of-thought pipelines [36] suggest that segmentation tools such as SAM can automatically generate semantic masks, enabling edited views without human annotation. Similarly, procedural domains like Mazes [19] offer precise rendering control, allowing exact synthesis of counterfactual views similar to ours. We leave a systematic exploration of such hybrid and multi-domain settings to future work.